The length of a rectangular field is three times its width.
The perimeter of the field is 144 cm. Find it's area

Answer by ewatrrr(24785)

 
Hi,

P = 2L + 2w = 144cm
L+ w = 72 L = 3w
3w + w = 72
4w = 72
w = 18cm and L = 54cm
A = Lw = 972cm^2